OpenStudy (anonymous):

I solved 2 cos(2x)+3=2 using inverse and then solved the same equation using the double angle formula, why are the answers different? what is the correct method?

OpenStudy (anonymous):

When I use the inverse I get 60 and 120 degrees When I use the cos (2x) double angle formula, I get 60 120 240 and 300 When I check with the TI83 I got 4 answers 60 120 240 and 300

OpenStudy (zehanz):

You could even find an infinite number of solutions. You use degrees. Does that imply your solutions have to be between 0 and 360?

OpenStudy (zehanz):

You found 2x = 120, or 2x = 240, so x = 60 or x = 120 (divide by 2). But: Every time you add 360, you've got a solution again, so the original solutions were actually: 2x = 120 + k*360, or 2x = 240 + k*360, where k is an integer. Again, to find x, divide by 2: x = 60 + k * 180 or x = 120 + k*180. If x has to stay in the interval [0, 360], you have solutions: 60, 120, 240, 300.